Solve each system by the substitution method (please):
2x - y = -5
x + 5y = 14
Thank you ^_^

Answer by oberobic(2304)

2x -y = -5
x +5y = 14
.
x = -5y +14
.
Substitute for 'x' in the other equation
.
2(-5y+14) -y = -5
-10y +28 -y = -5
-11y = -33
y = -33/-11
y = 3
.
x = -5y + 14
x = -5(3) + 14
x = -15 +14
x = -1
.
Answer: x = -1, y = 3
.
Check the answer
.
2x - y = -5 ?
2(-1) - 3 = -5
true
.
x + 5y = 14 ?
-1 +5(3) = 14
true
.
Done.